OpenStudy (anonymous):

anyone know how to Simplify this ??<3 (2 - 3i)2(2 + 3i)2

OpenStudy (anonymous):

(2-3i)^2(2+3i)^2 = (4 - 12i - 9) (4 + 12i - 9) = ((4-9)^2 - (12i)^2) = (-5)^2 - 12^2 * i^2 = 25 + 144 = 169

OpenStudy (anonymous):

or rearrange multipliers like this: (2-3i)(2+3i) * (2-3i)(2+3i) = (4 + 9) *( 4+ 9) = 13^2 = 169. It's simplier.